# Data-driven Virtual Test-bed of the Blown Powder Directed Energy Deposition Process

### v1

@@ 2.3 Uncertainty Quantification
Uncertainty quantification (UQ) can be approached from various perspectives [37]. In the context of this study, UQ was employed to ascertain the recorded sensor noise observed in the process. The assumptions inherent to dynamic mode decomposition imply that the model is devoid of the characteristic noise inherent to the process. This study specifically aimed to determine the noise frequency and amplitude. To achieve this, the OMM data was initially segregated based on the laser state, meaning that uncertainty quantification was only performed on data collected while the laser was active. The data was then further categorized by either scan rate ($mm/min$) or laser pulse length ($s$). These quantities aim to describe transient versus steady-state process behavior through slightly different means. With a fixed recording frequency, the scan rate ($mm/min$) directly affects the spatial resolution that can be captured, suggesting the potential existence of a varying noise component. In contrast, laser pulse length attempts to capture the ratio of transient to steady-state processing, with short pulses having a high ratio and long pulses a low ratio, effectively capturing the tension between these states.
After appropriately filtering the data set, the objective was to ascribe both a noise frequency and amplitude to the data. Broadly, frequency was characterized by the Fast Fourier Transform (FFT) of the signal, and amplitude was quantified as a signal-to-noise ratio (SNR). Specifically, we defined the noise frequency as the largest intensity peak of the FFT that is not the bias frequency, and the SNR as ($SNR=\mu/\sigma$). These definitions, while subject to refinement, were deemed applicable for this initial exploration.
Finally, both the noise frequency and amplitude were separately fitted to a function to facilitate the application of the UQ envelope to the DMDc state-space model output. For this purpose, a Gaussian Process (GP) model was employed due to its advantageous properties, including the provision of uncertainty bounds and ease of sampling [32]. GP models have been utilized in numerous AM related regression problems with varying levels of success, and we believe it is well-suited to the problem at hand [2, 29]. The primary drawback of GP models is their computational complexity, which is ($O(n^{3})$), imposing a practical limit on the number of data points that can be fit. To address this limitation, we employed multiple GP models to form a bagging regressor. Consequently, each GP model fit a small subset of the total data, and the final function represented an aggregation of all individual GP models.
@@ 3 Results and Discussion
3.1 Process Input Analysis
@@
Figure 2: Distributions of DED observables versus their uniform signal bounds, the so-called "intra-feature rank." Also indicated is the Wasserstein distance for each observable.
Figure 2 illustrates the distribution of each observable (melt pool size, temperature, and working distance) compared to a uniform distribution of the same size. Above each plot, the calculated Wasserstein distance is displayed, with the minimal distance for this metric being zero. A comprehensive table, including these observables as well as all other recorded signals, is provided in the supplemental material section. From the figure, it is evident that neither the test nor the training data is sufficiently diverse to approach a uniform distribution, yet both are full rank. Notably, the distributions of the test and training data are very similar, indicating that the trained model will predominantly perform interpolation rather than extrapolation.
3.2 DMDc Solutions
The model was trained using the DMDc algorithm on roughly 1.5 million data points. The training process took 3.76 $s$ followed by 21.96 $s$ to reproduce the data. The test data set contained 150 thousand data points which was simulated in 2.12 $s$. This was performed on a PC with an 11th Gen Intel®i7 @ 3.00GHz processor and 32GB of RAM. Figures 3a and 3b show the results of the reproductive and test case respectively for a single observable. Figures containing the results from the other observables are shown in Section 8.
Figure 3: DMDc results for training (a) and test data (b). (top) Time series of melt pool size for both ground truth and DMDc. Inset is the $R^{2}$ and mean squared error ($MSE$) values of the fit. (bottom) Residuals between the experimental data and model predictions.
Though the test data performs slightly better, in $R^{2}$ and $MSE$, than the training data, this difference is only slight and most likely within some margin of error, so the model seems to perform as well reproducing the training data as it did on the test data. The results are very promising as the melt pool size (mm) in Figure 3b is closely matched in the test case. What is left for section 8 are the results for the other observables where melt pool temperature (C) also performs very well, but working distance (mm) only achieves an $R^{2}$ in the low 70s. This could be due to a couple of reasons: (1) the physics of working distance cannot be captured by data-driven methods, and (2) the off state, where the camera receives no input, is represented by a $-1$ value which could be biasing the model.
3.3 Temporal Snapshot Frequency Analysis
@@
The experimental dataset was sub-sampled at various intervals to artificially create additional datasets at lower recording frequency. From these data sets, the entire developed DMDc pipeline was applied. The results are shown in Figure 4 with $R^{2}$ plotted as a function of hertz. With a simple spline used for illustration purposes, what is shown is the degradation of the model accuracy as a function of the coarseness of data collection. All three sensor observables indicate a downward trend in model accuracy below a given hertz value. For melt pool temperature this occurs below $50Hz$, and for both melt pool size and working distance it occurs below $20Hz$. The trend holds for all sensors even when there are marked differences in the total accuracy achievable. Applying this result spatially, at relevant process speeds (upper bound of $1500mm/min$), indicates a minimum of $1$ measurement every $1.25mm$ is required. For reference, the laser spot diameter is $1.2$ mm in the DED machine used for this study.
@@
3.4 Experimental Data Uncertainty Quantification
Looking at the UQ results presented in Figure 5, which are plotted as a function of laser pulse length, we observe several intriguing findings. As anticipated, the noise frequency decreases with increasing laser pulse length across all three observables: melt pool size, working distance, and melt pool temperature. This trend aligns with our expectations, as longer laser pulses typically result in more stable and consistent measurements, thereby reducing the noise frequency.
However, the SNR trends in reveal some unexpected patterns. For both melt pool size and working distance, the SNR behaves as expected, improving with longer laser pulse lengths. This improvement indicates that the measurements become more reliable and less affected by noise. In contrast, the SNR for melt pool temperature runs counter to this trend, showing a decrease with increasing laser pulse length. This anomaly suggests that other factors might be influencing the temperature measurements, potentially related to the thermal dynamics of the melt pool or the sensitivity of the temperature sensors.
Additionally, it is noteworthy that the working distance measurement generally performs poorly. This is evident from the high uncertainty in the GP model for noise and the very low SNRs observed. The high uncertainty indicates that the model struggles to accurately predict the noise characteristics for working distance, while the low SNRs suggest that the measurements are significantly affected by noise, making them less reliable.
The results are similar when examining the UQ results plotted with respect to scan speed as hown in Figure 6. For working distance measurements, we again see high noise variance and low SNRs, reinforcing the challenges in obtaining accurate and reliable measurements for this observable. For both melt pool size and temperature, the noise frequency appears to remain nearly constant across the range of scan speeds, indicating that scan speed does not significantly impact the noise characteristics for these observables.
However, the SNR plots for melt pool size and temperature exhibit some unexpected behavior. Specifically, for melt pool size, there seems to be a modeling artifact that exacerbates the underlying results. This artifact could be due to limitations in the GP model or other factors not accounted for in the analysis. When examining the furthest data points at either end of the scan speed range, we observe SNRs that align in magnitude with those of melt pool temperature. This alignment suggests that, despite the modeling artifact, the overall trend in SNR for melt pool size is consistent with that of melt pool temperature, highlighting the complex interplay between scan speed, noise, and measurement reliability.
Figure 5: Noise frequency (Hz) (top row) and amplitude (bottom row) assessment on the sensor states of the experimental training data. Frequency and Signal-to-Noise Ratio ($SNR$) given as a response to laser pulse length ($s$)
Figure 6: Noise frequency (Hz) (top row) and amplitude (bottom row) assessment on the sensor states of the experimental training data. Frequency and Signal-to-Noise Ratio ($SNR$) given as a response to scan rate ($mm/min$)
3.5 Combined Framework Predictions
@@
The output of this proposed framework combines the model predictions generated through the DMDc process and bounds those predictions using the UQ obtained from the experimental data. Providing expectation bounds gives machine operators real-time knowledge to understand when the process deviates from nominal conditions and requires intervention, either through parameter adjustments or by terminating the build if recovery is not feasible. These bounds also offer analysts a first-pass filtering mechanism to isolate anomalies. Linking OMM data streams to specific defects in non-single track contexts has proven challenging overall, though there have been some successes [41]. Analysis of the signal structure outside of these bounds may prove to be a viable strategy in making the problem more tractable.
Figure 7: Combined DMDc predictions with UQ applied as bounding envelope. Shown in (a) is melt pool size (mm) and melt pool temperature (C) in (b). Insets are 3D comparisons of the ground truth test data and the DMDc predictions.
The process to achieve the control plots (model predictions + UQ) shown in Figures 7a & 7b follows the outlined path. The model ingests process inputs of the build to be printed and yields predictions of the sensor states. These predictions are then sorted and grouped by either laser pulse length ($s$) or by scan rate ($mm/min$). For each grouping, a call is made to the appropriate SNR model, which produces a value. We then back-solve for the appropriate ($\sigma$), which requires an interpolation function on ($\mu$) to account for discretization mismatch. The model prediction is then augmented by ($\sigma$) as shown in Equation 11 to produce the bounds.
Signal-to-Noise Ratio, SNR $=\mu/\sigma\text{,}$ (11)
rearranging for $\text{rearranging for }\sigma\text{ yields, }\sigma$ $=\mu/\text{SNR}\text{.}$
[…]
There has been no mention of incorporating the noise frequency component of the UQ analysis. Simply put, while it is important information to have—such as verifying that noise frequency does not correlate with typical electrical frequencies and their harmonics—it does not translate as well when applied to model predictions. It was attempted to sample from those points at the rate suggested by the noise frequency model and only apply the SNR to that subset of sampled points. However, beyond trying to match the experimental signal, which it doesn’t do, there are no further analytical benefits to adding salt-and-pepper noise to the model predictions. It doesn’t aid the operator or the analyst in the context mentioned previously. What remains to be seen is if the noise frequency model can be used to filter the experimental data rather than add the noise into the DMDc model.
Analyzing model performance through the perspective of the parity plot provides higher-level insight into the model. Figure 8 shows such a plot for all three observables under consideration. Ignoring the $-1$ value in the working distance ($mm$) plot, which represents the state where the camera is making no observations, we can see that overall the model under-predicts the correct working distance. This observation correlates with the fact that the ($R^{2}$) value for this state performed significantly worse than for melt pool size or temperature. It isn’t entirely clear if the inclusion of this $-1$ value within the experimental data set is biasing the result to cause the under-prediction. The parity plot shows that for a relatively limited experimental measurement, the model predicts a fairly large range of values.
Regarding the other two measured states, the models are performing well in terms of metrics. However, there are regions within the parity plot that show a bifurcated behavior of either under-prediction or over-prediction. The DMDc algorithm, driven by least squares, properly balances this bifurcation, maintaining good overall quality on the metrics. Also, in those parity plots are islands with very good experimental data to prediction agreement. Continued assessment of the bifurcated regions may indicate that the model is loosely “understanding” the physics applicable to the problem, even if only through a black-box method. The only reason to imply this is that the bifurcation regions in both the melt pool size and melt pool temperature occur around phase change areas. For melt pool temperature, the material used was Inconel 625, which has a melting temperature in the $1290$ to $1350$°C range [9]. For melt pool size, the laser spot diameter is $1.2mm$, and therefore the concept of a melt pool doesn’t really make sense below a certain value. This could also mean that incorporating additional physics could lead to better accuracy of the model, even if a pathway to that incorporation is currently unknown.
Figure 8: Parity plots for all three states measured states. Highlighted in orange are areas where the model is predicting well. Highlighted in blue are areas where the model could perform better, however, due to symmetry in the parity plot overall error remains low. Lastly, working distance ($mm$) is almost entirely under-predicted by the DMDc model

### v2

@@ 2.4 Uncertainty Quantification
Uncertainty quantification (UQ) can be approached from various perspectives [30]. In this study, UQ was derived as a result of the LpOCV methodology employed to train the DMDC state-space model. The DMDc algorithm, which is based on least squares regression, tends to balance errors, resulting in normally distributed residuals. Through the cross-validation process of sampling, we have bootstrapped error bounds on the model, with the assumption that these bounds generalize to population bounds and can be utilized as control plots with confidence. To validate the UQ error bounds, we analyzed the frequency spectra of both the experimental training dataset and the resulting state-space model predictions for structural similarity. For this purpose, we constructed spectrograms of each observable, with the independent and dependent variables being laser pulse length ($s$) and frequency ($Hz$), respectively. Typically, spectrograms include time as the abscissa; however, our objective is to describe the transient versus steady-state process behavior. Therefore, we selected laser pulse length as the independent variable. Laser pulse length attempts to capture the ratio of transient to steady-state processing behavior, with short pulses exhibiting a high ratio and long pulses a lower ratio, effectively capturing the tension between these states.
To generate these spectrograms, the complete training dataset was analyzed by identifying the laser pulse lengths within the commanded laser power ($W$)feature. Commanded laser power is the input feature which details time points where the laser is active and at what intensity. When plotted, this feature appears as a series of square waves of varying lengths. Once the pulse lengths and their index locations were determined, the dataset was grouped according to these pulses, some of which contained multiple observations. A fast Fourier transform (FFT) was then applied to each individual pulse. If multiple observations existed within a grouping, the Fourier transforms were averaged to provide a representative spectrum for a given pulse length. Since these spectra are not necessarily of the same vector length, the data were linearly interpolated to fit onto a uniform rectangular grid, which could then be plotted for comparative analysis.
@@ 3 Results and Discussion
3.1 OMM Feature Selection and Curation
@@
Table 2 presents the results of the VIF feature selection algorithm. The algorithm removes one feature per iteration and eliminated a total of seven features. A commonality among most of the features removed, particularly in iterations 3 through 6, is that they are monotonically increasing. In iteration 1, a redundant column was removed because there is an existing feature called "Contour Flag (0/1)" that is exactly opposite. Notably, after iteration 1, the matrix rank and the dataset column space were equal; however, the VIF algorithm continued to identify collinear features. This observation serves as a loose verification that matrix rank is a necessary but not sufficient condition for feature selection. The final selected input features after the VIF algorithm as well as their distributions are summarized in Figure S1.
VIF Feature Selection Results
VIF Algorithm Feature Excluded Rank Indicator
Iteration 1 Infill Flag (0/1) Data Features = 28, Matrix rank = 27
Iteration 2 A Position (deg) Data Features = 27, Matrix rank = 27
Iteration 3 Current Layer (#) Data Features = 26, Matrix rank = 26
Iteration 4 Ticks Data Features = 25, Matrix rank = 25
Iteration 5 Part Distance Traveled (mm) Data Features = 24, Matrix rank = 24
Iteration 6 Program Time (s) Data Features = 23, Matrix rank = 23
Iteration 7 Enclosure 02 (ppm) Data Features = 22, Matrix rank = 22
Table 2: Results of the feature selection algorithm. Column 1 indicates the iteration number, column 2 designates the feature ultimately excluded from the final dataset, and column 3 displays the number of columns in the current dataset as of that iteration and the corresponding matrix rank.
With the inter-signal analysis completed and features set, the next step is to proceed with the intra-signal analysis of those features. As described in the section 2, the Wasserstein distance was employed to quantify the distance between distributions, noting that a Wasserstein distance of zero indicates identical distributions. Table 3 presents these results within the context of the LpOCV framework. Generally, for all three observables, there is significantly more distributional distance between test/train dataset and the uniform distribution. Conversely, the test and train distributions exhibit a lower Wasserstein distance. As shown in Figure 2 for a random data split, the test and train datasets are distributionally similar, indicating that a derived state-space model will primarily perform interpolation rather than extrapolation.
Wasserstein Distance Results (LpOCV)
Melt Pool Size (mm) Melt Pool Temperature (C) Realtime Working Distance (mm)
Mean 95% CI Mean 95% CI Mean 95% CI
Train→Uniform 1.375 $\pm$.062 Train→Uniform 295.887 $\pm$16.564 Train→Uniform 5.152 $\pm$.119
Test→Uniform 1.378 $\pm$.585 Test→Uniform 309.569 $\pm$121.963 Test→Uniform 5.4 $\pm$1.007
Test→Train .257 $\pm$.422 Test→Train 105.524 $\pm$175.706 Test→Train .525 $\pm$.652
Table 3: Table of Wasserstein distance results for the three states of interest. Results compare both the test and train datasets to a uniform distribution and lastly compares the test to the training datasets all using the LpOCV methodology. 95% confidence intervals provided as a measure of variance in the received results.
Other points to highlight in Table 3 include that both test and train datasets, when compared to the uniform distribution, generally have relatively the same Wasserstein distance, while the variance decreases as the data size increases (i.e., the variance for the test versus uniform is higher than for the train versus uniform). Notably, there is high variance in the Wasserstein distance reported for melt pool temperature. The origins and effects of such a large variance are currently unknown. Additionally, the impact of a large Wasserstein distance on the final model is not fully understood and will be an area of focus for future research. Lastly, Figure S1 provides box-and-whisker plots of all other process inputs, serving as another visualization tool for assessing distributional overlap.
@@
Figure 2: Distributions of DED observables versus their uniform signal bounds.
3.2 State-Space Model: DMDc Results
The state-space, surrogate model was trained using the DMDc algorithm on approximately 1.5 million data points at a rate of $2.5\mu s$ per data point, and achieving an inference rate of $14.35\mu s$ per data point on relatively standard hardware (11th Gen Intel®i7 @ 3.00GHz). The training results can be observed in Table 4
Table 3 lists the sensor observables—melt pool size ($mm$), melt pool temperature ($C$), and real-time working distance ($mm$)—evaluated using the metrics $R^{2}$ and root mean squared error ($RMSE$) for both the reproductive and test cases. The $R^{2}$ results indicate a common trend where the training case outperforms the test case which is to be expected. The performance of the state-space model for melt pool size and melt pool temperature is similar between the training and test evaluations and also enjoys a high percentage of data explained by the model. Across all observables, there is an observed increase in variance for the test case. For melt pool size and temperature, the error is considered reasonable, with the melt pool size showing an average error of about 280 microns and melt pool temperature approximately 140°C. This corresponds to a 23% error in melt pool size relative to the laser beam diameter and a 10% error relative to the melting temperature of Inconel 625. It remains to be determined what constitutes an acceptable level of error for these quantities, given that the underlying model explains a high percentage of the OMM data. Nonetheless, minimizing error is preferable for improving model accuracy and reliability.
DMDc Results (LpOCV)
Melt Pool Size (mm) Melt Pool Temperature (C) Realtime Working Distance (mm)
Mean 95% CI Mean 95% CI Mean 95% CI
$R^{2}$ (Train) .96 $\pm$.01 $R^{2}$ (Train) .96 $\pm$.00 $R^{2}$ (Train) .72 $\pm$.02
$R^{2}$ (Test) .94 $\pm$.04 $R^{2}$ (Test) .95 $\pm$.02 $R^{2}$ (Test) .64 $\pm$.25
RMSE (Train) .25 $\pm$.02 RMSE (Train) 138.6 $\pm$3.0 RMSE (Train) 2.26 $\pm$.10
RMSE (Test) .28 $\pm$.12 RMSE (Test) 141.6 $\pm$36.0 RMSE (Test) 2.42 $\pm$.93
Table 4: Table of $R^{2}$ and $RMSE$ results for the three states of interest. Results reported for performance on both the test and train datasets. All DMDc results obtained using the LpOCV methodology and 95% confidence intervals provided as a measure of variance in the received results.
The successes of the state-space model for melt pool size and temperature are unfortunately contrasted by its relatively poor performance on real-time working distance. The model explains roughly 64% of the data, with an average error of nearly 2.5 ($mm$). While the errors for melt pool size and temperature are reasonable, an error of this magnitude for most DED machines can lead to significant discrepancies in powder delivery to the melt pool. There are a couple of potential reasons for the model’s poor performance:
  • The physics of working distance may not be sufficiently captured by data-driven methods, necessitating a robust physics-based, white-box approach.
  • Equipment peculiarities, particularly when the camera cannot observe the melt pool, may affect performance. This can occur when the laser is off or when the laser is on but the melt pool is, for various reasons, out of the single-perspective camera’s view. During these "off" states, the recorded value is -1, which, while technically nonsensical, could introduce bias into the model.
We believe there is no compelling reason why the physics of working distance cannot be captured by data-driven methods. Therefore, in the next section, we will address the potential bias introduced by the camera’s off state and explore strategies to mitigate its impact on model performance.
3.2.1 Addressing Working Distance: Imputation Study
To address the potential bias introduced by the camera’s off state, we have chosen to perform an imputation study. In this study, all camera off state values occurring while the laser is firing are imputed to a non-negative value. Although imputation is typically used to handle missing data, we consider the camera off state as equivalent to missing data in this context. There are various strategies for imputing data, such as substituting with the mean (or some other statistic) of the feature or using a constant value. We have opted for linear interpolation to impute the camera off state. In this approach, for any subset requiring imputation, the nearest values outside of that subset are used to form a line. The interpolated values on this line then serve as replacements for the missing data.
DMDc Imputation Study Results (LpOCV)
Melt Pool Size (mm) Melt Pool Temperature (C) Realtime Working Distance (mm)
Original→New Original→New Original→New
Mean 95% CI ($\pm$) Mean 95% CI ($\pm$) Mean 95% CI ($\pm$)
$R^{2}$ (Train) .96→.95 .01→.01 .96→.96 .$0$72→.88 .02→.01
$R^{2}$ (Test) .94→.95 .04→.03 .95→.96 .02→.02 .64→.88 .25→.04
RMSE (Train) .25→.26 .02→.01 138.6→138.1 3.0→5.2 2.26→1.48 .10→.06
RMSE (Test) .28→.27 .12→.06 141.6→140.5 36.0→28.0 2.48→1.45 .93→.28
Table 5: Table of $R^{2}$ and $RMSE$ results of the imputation study for the three states of interest. Results reported for performance on both the Test and Train datasets with green and red highlighting indicating a performance improvement or deterioration, respectively. All DMDc results obtained using the LpOCV methodology and 95% confidence intervals provided as a measure of variance in the received results.
After performing imputation on the realtime working distance feature, we conducted a complete LpOCV training of the state-space model using the DMDc approach. The results, presented in Table 5, show a marked improvement in the amount of explained variance for realtime working distance, along with a reduction in the overall error for this observable. Table 5 also highlights improvements in various metrics across all three observables. Since the state-space model is a coupled system, the enhancement in one observable positively influenced the performance of the other two observables. These improvements are evident in the increased explained variance, decreased average error, and reduced model variance. This promising outcome supports the notion that realtime working distance can indeed be effectively captured using data-driven methods. The improvements across multiple metrics underscore the potential of this approach in enhancing the accuracy and reliability of the model.
3.2.2 OMM Recording Frequency Analysis
@@
The experimental dataset was sub-sampled at various intervals to artificially create additional datasets at lower recording frequency. From these data sets, the entire developed DMDc pipeline was applied. The results are shown in Figure 3 with $R^{2}$ plotted as a function of hertz. With a simple spline used for illustration purposes, what is shown is the degradation of the model accuracy as a function of the coarseness of data collection. All three sensor observables indicate a downward trend in model accuracy below a given hertz value. For melt pool temperature this occurs below 50 Hz, and for both melt pool size and working distance it occurs below 20 Hz. Applying this result spatially, at relevant process speeds (upper bound of 1500 mm/min), indicates a minimum of 1 measurement every 1.25 mm is required. For reference, the laser spot diameter is 1.2 mm in the DED machine used for this study.
@@
3.3 Uncertainty Validation
As described in section 2, we constructed spectrograms of each observable, with the independent and dependent variables being laser pulse length ($s$) and frequency ($Hz$), respectively. Upon reviewing the results shown in Figure 4a and 4b, all spectrograms demonstrate peak intensities that decay with increasing laser pulse length, as previously hypothesized. This suggests that the noise component is dependent on the ratio of transient to steady-state deposition. Notably, none of the plots exhibit peak intensities above 3 $Hz$. Consequently, we have cropped these plots to display a maximum of 1 $Hz$ to enhance the visibility of the spectrogram structure. Additionally, all intensity peaks have been normalized to a range of 0 to 1. Given a Nyquist frequency of 50 $Hz$, it is possible that we are missing electrical noise and other higher frequency sources. The final result shows structural parity between the experimental data and model data spectrograms, indicating that the model accurately captures the key features of the experimental observations.
Figure 4: Experimental data spectrograms (left) and state-space model spectrograms (right) for all three observables. These spectrograms are frequency ($Hz$) as a function of laser pulse length ($s$).
3.4 State-Space Model + Uncertainty Quantification Predictions
@@
The output of this proposed framework combines the state-space model predictions generated through the DMDc training process and bounds those predictions using the UQ obtained from the LpOCV. Providing expectation bounds gives machine operators realtime knowledge to understand when the process deviates from nominal conditions and requires intervention, either through parameter adjustments or by terminating the build if recovery is not feasible. These bounds also offer analysts a first-pass filtering mechanism to isolate anomalies. Linking OMM data streams to specific defects in non-single track contexts has proven challenging overall, though there have been some successes [35]. Analysis of the signal structure outside of these bounds may prove to be a viable strategy in making the problem more tractable.
Figure 5a and 5b presents the combined results for melt pool size in both the training and test cases. The maximum and minimum error bounds are included, with portions of the experimental data that fall outside these bounds highlighted in red. Corresponding figures for other observables are available in the appendix (Figures S2a, S2b, and S3a, S3b). The error bounds are calculated by taking the model prediction $\pm$ the $RMSE$$\pm$ the 95% confidence interval. On the right-hand side of the figures, histograms of the residuals for both the training and test cases are displayed. The training residuals are normally distributed, while the test case residuals exhibit a small amount of skewness. This indicates that while the model performs well on the training data, there is some deviation in the test data that may need further investigation or model refinement.
Figure 5: DMDc state-space model results for reproductive, or training, (a) and test data (b). (Left) Time series of melt pool size for both ground truth and DMDc state-space model. (Right) Histogram of residuals between the experimental data and model predictions.
A parity plot (Figure 6) provides a higher-level perspective on model performance for the three observables under consideration. All three measured states perform well according to the metrics. Additionally, a separate linear regression model (dashed line) is plotted to indicate areas of under- or over-prediction. For all three observables, there is over-prediction at lower values, which transitions to under-prediction at higher values. Notably, real-time working distance experiences the most significant under-prediction at process-relevant distances. Interestingly, bifurcated behavior is observed in both melt pool size and temperature. This suggests that the model may be loosely capturing the underlying physics of the problem, albeit through a data-driven, black-box approach. The bifurcation regions in both melt pool size and temperature occur around phase change areas. For melt pool temperature, the material used is Inconel 625, which has a melting temperature in the range of 1290 to 1350°C [8]. Regarding melt pool size, the laser spot diameter is 1.2 mm, implying that the concept of a melt pool becomes less meaningful below a certain size. This observation suggests that incorporating additional physics into the model could enhance its accuracy, even though the pathway for such incorporation is currently unknown. Further investigation into these bifurcated regions could provide insights into improving the model’s understanding of the relevant physical phenomena.
Figure 6: Parity plots for all three measured states are presented.